A hiker goes six miles east and then turns south. If the hiker finishes 7.2 miles from the starting point, how far did the hiker go?

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

a^2+b^2=c^2
a=6
c=72
6^2+b^2=7.2^2
b^2+36 = 51.84
b = 3.97995

Answer by stanbon(75887)

A hiker goes six miles east and then turns south. If the hiker finishes 7.2 miles from the starting point, how far did the hiker go?
-----------
Use Pythagoras:
6^2 + s^2 = 7.2^2
-------
s^2 = 15.84
-------
south = 3.98 miles
=================
Total walk:: 6 +3.98 = 9.98 miles
==================
Cheers,
Stan H.
=================